OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1

OpenStudy (anonymous):

0 solutions

OpenStudy (anonymous):

they are the same thing oh so maybe infinite...

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

well it is infinite they are the same :P

OpenStudy (anonymous):

X+Y =1 Or, Y=-X +1 Which is same as, X+Y=1 So, In both cases, Y=-X +1 So equating both Y and Y, Y=Y -X+1=-X+1 -X=-X X=X Infinitely many solutions because X=X.